40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Entry  Wed Dec 4 19:00:37 2019, Noisemon board modified, DailyProgress, , Noisemon board modifications 
    Reply  Thu Dec 5 12:37:58 2019, Noisemon board modified, DailyProgress, , Noisemon board modifications ModificationComparison.pdf
       Reply  Thu Dec 5 20:42:55 2019, Duo, DailyProgress, Electronics, Noisemon board modifications NoDrive.pdf10HzSine.pdfTF.pdf
Message ID: 1837     Entry time: Thu Dec 5 20:42:55 2019     In reply to: 1836
Author: Duo 
Type: DailyProgress 
Category: Electronics 
Subject: Noisemon board modifications 

Fully board successful with modification. The saturation issue has been fixed on all the four channels.

Attachment 1: NoDrive.pdf  40 kB  | Hide | Hide all
NoDrive.pdf
Attachment 2: 10HzSine.pdf  39 kB  | Hide | Hide all
10HzSine.pdf
Attachment 3: TF.pdf  22 kB  | Hide | Hide all
TF.pdf
ELOG V3.1.3-